OpenStudy (anonymous):

graph the equation and plot three points 2y=3x+11

OpenStudy (anonymous):

this is equation of the line. to plot points - identify 3 random x, plug-in in equation & find corresponding Y

OpenStudy (anonymous):

for a example: x=0, then y=11/2 right?

OpenStudy (anonymous):

if x=-1, then y=8/2 =4 if x=-3, then y=2/2=1 right?

OpenStudy (anonymous):

so, you have 3 coordinates: (0,11/2), (-1,4), (-3,1) put it /identify on coordinate plane